# LLM4GEN: Leveraging Semantic Representation of LLMs for Text-to-Image Generation

### v1

@@ Abstract.
Diffusion Models have exhibited substantial success in text-to-image generation. However, they often encounter challenges when dealing with complex and dense prompts that involve multiple objects, attribute binding, and long descriptions. This paper proposes a framework called LLM4GEN, which enhances the semantic understanding ability of text-to-image diffusion models by leveraging the semantic representation of Large Language Models (LLMs). Through a specially designed Cross-Adapter Module (CAM) that combines the original text features of text-to-image models with LLM features, LLM4GEN can be easily incorporated into various diffusion models as a plug-and-play component and enhances text-to-image generation. Additionally, to facilitate the complex and dense prompts semantic understanding, we develop a LAION-refined dataset, consisting of 1 million (M) text-image pairs with improved image descriptions. We also introduce DensePrompts which contains 7,000 dense prompts to provide a comprehensive evaluation for the text-to-image generation task. With just 10% of the training data required by recent ELLA, LLM4GEN significantly improves the semantic alignment of SD1.5 and SDXL, demonstrating increases of 7.69% and 9.60% in color on T2I-CompBench, respectively. The extensive experiments on DensePrompts also demonstrate that LLM4GEN surpasses existing state-of-the-art models in terms of sample quality, image-text alignment, and human evaluation. The project website is at: https://xiaobul.github.io/LLM4GEN/
Text-to-Image Diffusion Models, Large Language Models, Text-Image Alignment
[…]
Figure 1. Image generation comparison using short and dense prompts across SDXL (Podell et al., 2023), Playground v2 (Li et al., [n. d.]), PixArt-$\alpha$ (Chen et al., 2023b), and our proposed LLM4GENSDXL. The colored text denotes critical entities or attributes.
1. Introduction
@@
Recently, diffusion models (Sohl-Dickstein et al., 2015; Dhariwal and Nichol, 2021; Song et al., 2020b; Yang et al., 2023; Chen et al., 2023a; Huang et al., 2022) have made significant progress in text-to-image generation models, such as Imagen (Saharia et al., 2022), DALL-E 2/3 (Ramesh et al., 2022; Betker et al., 2023), and Stable Diffusion (Rombach et al., 2022; Podell et al., 2023). However, they often encounter challenges in generating images given complex and dense prompt descriptions, such as attribute binding, orientation descriptions, and multiple objects. These limitations may stem from restrictions within the parameters and structure of the text encoder (Huang et al., 2023).
With the emergence of powerful linguistic representations from Large Language Models (LLMs), there has been an increasing trend in leveraging LLMs to aid in text-to-image generation. Current methods mainly consist of two categories: LLM-guidance models (Yang et al., 2024; Feng et al., 2022) and LLM-alignment models (Saharia et al., 2022; Wu et al., 2023b; Hu et al., 2024). LLM-guidance models aim to harness the robust encoding and decoding capabilities of LLMs and external models, such as the Layout model, to enhance the representation of text embeddings, subsequently generating images, as illustrated in Fig. 2 (a). However, these methods necessitate the separation of LLMs and external models from the text-to-image models, resulting in redundancy in both inference time and framework. While LLM-alignment models utilize LLMs instead of the vanilla text encoder to capitalize on their superior representational power. This strategy requires a substantial amount of training data to align the representation of LLMs with the diffusion model, imposing a redundant cost on training resources and time, as depicted in Fig. 2 (b).
@@
Figure 2. Architecture comparison between (a) LLM-guidance models(Lian et al., 2024; Yang et al., 2024), (b) LLM-alignment models (Wu et al., 2023b; Hu et al., 2024), and (c) our proposed LLM4GEN.
@@
To tackle the aforementioned challenges, we propose an end-to-end text-to-image framework named LLM4GEN, implicitly harnessing the powerful semantic representations of LLMs to bolster the representation of the original text encoder for text-to-image GENeration, as depicted in Fig. 2 (c). Specifically, we design an efficient Cross-Adapter Module (CAM) to implicitly integrate the semantic representation of LLMs with original text encoders that have limited representational capabilities, such as CLIP text encoder (Radford et al., 2021). We apply cross-attention on the representation of both encoder-only and decoder-only LLMs, e.g. Llama (Zhang et al., 2024) and T5 (Raffel et al., 2020), alongside CLIP text embedding, and then concatenate the fused text embedding with CLIP text embedding. The utilization of our CAM substantially improves the performance of text-to-image diffusion models, regardless of whether encoder-only or decoder-only LLMs are used. Furthermore, it preserves the representations produced by original text encoders, thereby diminishing the requirement for extensive training data. Our designed LLM4GEN can be seamlessly integrated into existing diffusion models like SD1.5 (Rombach et al., 2022) and SDXL (Podell et al., 2023). As evidenced in Fig. 1, our proposed method exhibits strong performance in text-to-image generation.
@@
To comprehensively assess the image generation capabilities of text-to-image models, we develop a comprehensive benchmark named DensePrompts, an extension of T2I-CompBench (Huang et al., 2023), which incorporates over 7,000 compositional prompts. The construction of this benchmark involves leveraging LLMs for complex text descriptions, followed by manual refinement. Results from performance metrics and human evaluations consistently demonstrate that LLM4GEN’s representational capability surpasses other existing methods. Additionally, we introduce a LAION-refined dataset, which is a subset of the open-source LAION (Schuhmann et al., 2021) dataset, comprising 1M text-image pairs with refined image descriptions. Both the DensePrompts and LAION-refined dataset will be available to the community for further research.
@@
  • We propose an end-to-end text-to-image framework called LLM4GEN to enhance the text-to-image alignment of diffusion models.
  • We introduce a simple but efficient Cross-Adapter Module that integrates robust representation of LLMs into original text encoders that have limited representational capabilities, thereby enhancing semantic understanding for diffusion models.
  • We introduce DensePrompts, a comprehensive evaluation benchmark for text-to-image generation, and LAION-refined dataset, a meticulously curated training dataset.
@@
  • Experiments show that LLM4GEN exhibits superior performance in sample quality, image-text alignment, and human evaluation compared with existing state-of-the-art models.
2. Related Work
2.1. Large Language Models
@@
Large language models (LLMs) (Chang et al., 2023) have shown powerful generalization ability in various NLP tasks, e.g., text generation, question answering. LLMs are built on the transformer (Vaswani et al., 2017) architecture and guided by the scaling law. Recent LLMs, e.g., GPTs (Brown et al., 2020), LLaMA (Touvron et al., 2023), OPT (Zhang et al., 2022), BLOOM (Scao et al., 2022), GLM (Zeng et al., 2022), PaLM (Chowdhery et al., 2022) are all equipped with billions of parameters, enabling the intriguing capability for in-context learning and demonstrating excellent zero-shot performance across various tasks. Certain Multi-modal LLMs (MLLMs) (Achiam et al., 2023; Zhu et al., 2023; Bai et al., 2023; Wu et al., 2023a) have effectively integrated LLMs with other modalities, e.g., visual and audio, facilitating more intelligent interactions through the assistance of LLMs. BLIPs (Li et al., 2022, 2023a), LLaVA (Liu et al., 2023) enhance the synergy of visual understanding and language processing by projecting the visual output to the input layer of LLMs. (Pang et al., 2024) shows that the frozen LLMs can further integrate visual understanding. For the text-to-image generation, recent works (Lian et al., 2024; Koh et al., 2023; Sun et al., 2024; Yang et al., 2024) utilize LLMs to generate the refined text prompts or bounding box layouts (Li et al., 2023b) for high-quality image generation. However, these existing works only consider LLMs as a simple condition generator, e.g., text prompts or layout planning. In this paper, we harness the representation capabilities of LLMs to enhance text-to-image generation, underscoring the significance of their representational power beyond simple text output.
Figure 3. The overview of LLM4GEN. (a) The inference pipeline. (b) Cross-Adapter Module.
2.2. Text-to-Image Diffusion Models
@@
Text-to-image generation aims to create images based on given textual descriptions. The diffusion-based models (Sohl-Dickstein et al., 2015; Song and Ermon, 2019; Ho et al., 2020; Song and Ermon, 2020; Song et al., 2020b) have demonstrated remarkable performance in image generation, providing enhanced stability and controllability. These models employ a forward process by adding Gaussian noise to input images and can generate high-quality images with intricate details and diversity through an inverse process from random Gaussian noise. GLIDE (Nichol et al., 2022) and Imagen (Saharia et al., 2022) utilize CLIP (Radford et al., 2021) text encoder to enhance the image-text alignment. Latent Diffusion Models (LDMs) (Rombach et al., 2022) have been proposed to transfer the diffusion process from pixel to latent space, improving efficiency and image quality. Recent models such as SD-XL (Podell et al., 2023), DALL-E 3 (Betker et al., 2023), and Dreambooth (Ruiz et al., 2023) have significantly enhanced image quality and text-image alignment using various perspectives, such as training strategies and scaling training data. Despite these notable advancements, generating high-fidelity images aligned with complex and dense textual prompts remains challenging. In this paper, we propose LLM4GEN that implicitly leverages the robust representation capabilities of LLMs to facilitate image generation from textual descriptions.
3. Method
3.1. Preliminaries
Diffusion models convert standard Gaussian noise into realistic images through two processes: the forward process and the reverse process (Ho et al., 2020). In the forward process, Gaussian noise is gradually added to the data $x_{0}\sim q(x_{0})$ to obtain a Gaussian distribution with $T$ steps. The reverse process aims to generate the data from an initial Gaussian noise as follows (Ho et al., 2020):
(1) $p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\Sigma_{\theta}(x_{t},t))$
where $\mu_{\theta}$ can be obtained by predicting the Gaussian noise $z(x_{t},t)$ as follows:(Ho et al., 2020)
(2) $\mu_{0}(x_{t}|t)=\frac{1}{\sqrt{\overline{a}_{t}}}(x_{t}-\frac{\beta_{t}}{\sqrt{1-\overline{a}_{t}}}\epsilon_{\theta}(x_{t},t))$
where $\epsilon(x_{t})$ is the noise added by the forward process predicted by the neural network.
Conditional diffusion models(Ho and Salimans, 2022) can be guided in classifier-free guidance, which replaces the explicit classifier with an implicit classifier that does not require calculating the explicit classifier and its gradient. By using classifier-free guidance, the predicted noise can be described as(Ho et al., 2020):
(3) $\overline{\epsilon}_{\theta}(x_{t},t,y)=(\omega+1)\epsilon_{\theta}(x_{t},t,y)-\omega\epsilon_{\theta}(x_{t},t)$
where $\omega$ is the guidance scale,$y$ is the condition added to guide the diffusion model.
Our work utilizes latent diffusion models (LDMs)(Rombach et al., 2022), consisting of three components: a text encoder such as CLIP(Radford et al., 2021) for extracting text embeddings, a variational autoencoder (VAE)(Van Den Oord et al., 2017) with an encoder $\mathcal{E}$ to encode images into a low-dimensional latent space, and a decoder $\mathcal{D}$ to reconstruct images from the encoded latent vectors, and a UNet for predicting noise during the diffusion process. The encoder $\mathcal{E}$ maps the input image to the latent space $z_{t}$, which is efficient and low-dimensional. The underlying UNet(Ronneberger et al., 2015) is constructed using 2D convolutional layers and reweighted bounds to focus on the most perceptually relevant features, as follows:
@@ 3.1. Preliminaries
(4) $L_{LDM}:=\mathbb{E}_{\varepsilon(x),\epsilon\sim\mathcal{N}(0,1),t}[\parallel\epsilon-\epsilon_{\theta}(z_{t},t)\parallel]_{2}^{2}]$
@@
3.2. LLM4GEN
3.2.1. Framework
The proposed LLM4GEN, which contains a Cross-Adapter Module (CAM) and the UNet, is illustrated in Fig. 3 (a). In this paper, we explore stable diffusion (Rombach et al., 2022; Podell et al., 2023) as the base text-to-image diffusion model, and the vanilla text encoder is from CLIP (Radford et al., 2021). LLM4GEN leverages the strong capability of LLMs to assist in text-to-image generation. The CAM extracts the representation of a given prompt via the combination of LLM and CLIP text encoder. The fused text embedding is enhanced by leveraging the pre-trained knowledge of LLMs through the simple yet effective CAM. By feeding the fused text embedding, LLM4GEN iteratively denoises the latent vectors with the UNet and decodes the final vector into an image with the VAE.
3.2.2. Cross-Adapter Module
(5) $Q=W_{q}(c_{l}),K=W_{k}(c_{t}),V=W_{v}(c_{t})$
(6) $c_{l}^{\prime}=\operatorname{CrossAttention}(Q,K,V)=\operatorname{softmax}\left(Q\cdot K^{T}\right)\cdot V$
(7) $c_{t}^{\prime}=\text{Concat}(\lambda*c_{l}^{\prime},c_{t})$
where the Concat denotes concatenation in the sequence dimension, and $\lambda$ is the balance factor.
Algorithm 1 LLM4GEN Pipeline
1:  Input: Pretrained UNet $\epsilon$, pre-trained text encoder $T_{\phi}$, pre-trained LLM $T_{L}$, Cross-Adapter Module $M$, LLaVA-7B $A$, training image-text pairs $\mathbb{S}=\{\mathbb{I},\mathbb{P}\}$
2:  Offline Process: Apply $A$ to enrich the captions of $\mathrm{S}$, get $\hat{\mathbb{P}}$ = $A$($\mathrm{I}$) to replace the original $\mathbb{P}$.
3:  $\rightarrow$ Begin Training.
4:  Freeze $T_{\phi}$ and $T_{L}$, training $M$ and $\epsilon$.
5:  while ${L_{simple}}$ not converged do
6:     Sample $p$ from $\mathbb{P}$; $t=T$
7:     $\mathbb{c_{t}^{\prime}}=M(T_{\phi}(p),T_{L}(p))$ using Eq. 7.
8:     Calculate ${L_{simple}}$ using Eq. 9.
9:     Backward ${L_{simple}}$ and update $\epsilon$, $M$.
10:  end while
11:  $\rightarrow$ End Training.
Then, we feed $c_{t}^{\prime}$ into the UNet:
(8) $x$ $=\operatorname{CA}(x,c_{t}^{\prime})=\lambda*\operatorname{CA}(x,c_{l})+\operatorname{CA}(x,c_{t})$
where $x$ denotes the latent noise, CA is the cross-attention module within the UNet module, which receives $z$ as the query and $c_{t}^{\prime}$ as the key and value.
Overall, our designed Cross-Adapter Module implicitly facilitates the strong representation of LLMs with a residual fusion manner, without utilizing extensive training data and resources to condition the latent vectors on text embeddings. Notably, our LLM4GEN is compatible with both decoder-only and encoder-only LLMs and we evaluate on Llama-2 7B/13B (Touvron et al., 2023) and T5-XL (Brown et al., 2020) in further experiments.
3.2.3. Training Objectives
Based on the framework described above, the training loss of LLM4GEN is formulated as:
(9) $L_{\text{simple }}=\mathbb{E}_{\boldsymbol{x}_{0},\boldsymbol{\epsilon},\boldsymbol{c}_{t},\boldsymbol{c}_{l},t}\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}\left(\boldsymbol{x}_{t},\boldsymbol{c}_{t},\boldsymbol{c}_{l},t\right)\right\|^{2}$
Besides, we randomly drop text embedding conditions in the training stage to enable classifier-free guidance in the inference stage:
(10) $\hat{\boldsymbol{\epsilon}}_{\theta}\left(\boldsymbol{x}_{t},\boldsymbol{c}_{t},\boldsymbol{c}_{l},t\right)=\omega\boldsymbol{\epsilon}_{\theta}\left(\boldsymbol{x}_{t},\boldsymbol{c}_{t},\boldsymbol{c}_{l},t\right)+(1-\omega)\boldsymbol{\epsilon}_{\theta}\left(\boldsymbol{x}_{t},t\right)$
The proposed LLM4GEN is further illustrated in Algorithm 2.
Figure 4. Pipeline of data construction. (a) The construction of DensePrompts benchmark. (b) The construction of the training dataset. We use GPT-4 (Achiam et al., 2023) and LLaVA (Liu et al., 2023) as the caption models, respectively.
Figure 5. Statistic of DensePrompts benchmark.
4. Dataset Construction
4.1. DensePrompts Benchmark
@@
A comprehensive benchmark is crucial for evaluating the image-text alignment of generated images. Current benchmarks, e.g., MSCOCO (Lin et al., 2014) and T2I-CompBench (Huang et al., 2023), primarily consist of concise textual descriptions, are not comprehensive enough to describe a diverse range of objects. Thus, we introduce a new comprehensive and complicated benchmark called DensePrompts, comprising lengthy textual descriptions.
@@
Initially, we collect 100 images from the Internet, comprising 50 real and 50 generated images, each with intricate details. Leveraging the robust image comprehension capabilities of GPT-4V (OpenAI, 2023), we utilize it to provide detailed descriptions for these 100 images, encompassing object attributes and their relationships, thereby generating comprehensive prompts abundant in semantic details. As depicted in Fig. 4 (a), we employ GPT-4 (Achiam et al., 2023) to produce massive long texts based on generated prompts mentioned above. DensePrompts provides more than 7,000 extensive prompts whose average word length is more than 40. Word statistics of DensePrompts are outlined in Fig. 5. To assess the performance, DensePrompts benchmark incorporates CLIP Score (Radford et al., 2021) and Aesthetic Score (Schuhmann, [n. d.]). Combining our proposed DensePrompts with T2I-CompBench, we establish a comprehensive evaluation in text-to-image generation.
Table 1. Evaluation results (%) on T2I-CompBench (Huang et al., 2023). The higher is better, and the best results are highlighted in bold.
@@
Composable Diffusion (Liu et al., 2022) 40.63 32.99 36.45 8.00 29.80 28.98
Structured Diffusion (Feng et al., 2022) 49.90 42.18 49.00 13.86 31.11 33.55
Attn-Exct v2 (Chefer et al., 2023) 64.00 45.17 59.63 14.55 31.09 34.01
GORS (Huang et al., 2023) 66.03 47.85 62.87 18.15 31.93 33.28
DALL-E 2 (Ramesh et al., 2022) 57.50 54.64 63.74 12.83 30.43 36.96
PixArt-$\alpha$ (Chen et al., 2023b) 68.86 55.82 70.44 20.82 31.79 41.17
ELLASDXL (Hu et al., 2024) 72.60 56.34 66.86 22.14 30.69 -
SD1.5 (Rombach et al., 2022) 37.65 35.76 41.56 12.46 30.79 30.80
LLM4GENSD1.5 45.34 43.28 51.52 13.12 30.94 32.33
$\Delta$ (Margin) +7.69 +7.52 +9.94 +0.66 +0.15 +1.53
SDXL (Podell et al., 2023) 63.69 54.08 56.37 20.32 31.10 40.91
LLM4GENSDXL 73.29 57.34 67.86 22.59 31.94 41.23
$\Delta$ (Margin) +9.60 +3.26 +11.52 +2.27 +0.84 +0.33
Table 2. Quantitative comparison on text-to-image generation models on the subset of MSCOCO (Lin et al., 2014) dataset.

### v2

@@ Abstract
Diffusion models have exhibited substantial success in text-to-image generation. However, they often encounter challenges when dealing with complex and dense prompts involving multiple objects, attribute binding, and long descriptions. In this paper, we propose a novel framework called LLM4GEN, which enhances the semantic understanding of text-to-image diffusion models by leveraging the representation of Large Language Models (LLMs). It can be seamlessly incorporated into various diffusion models as a plug-and-play component. A specially designed Cross-Adapter Module (CAM) integrates the original text features of text-to-image models with LLM features, thereby enhancing text-to-image generation. Additionally, to facilitate and correct entity-attribute relationships in text prompts, we develop an entity-guided regularization loss to further improve generation performance. We also introduce DensePrompts, which contains $7,000$ dense prompts to provide a comprehensive evaluation for the text-to-image generation task. Experiments indicate that LLM4GEN significantly improves the semantic alignment of SD1.5 and SDXL, demonstrating increases of 9.69% and 12.90% in color on T2I-CompBench, respectively. Moreover, it surpasses existing models in terms of sample quality, image-text alignment, and human evaluation.
Introduction
@@
Recently, diffusion models (Song et al. 2020; Rombach et al. 2022; Chen et al. 2023a) have made significant progress in text-to-image (T2I) generation models, such as Imagen (Saharia et al. 2022), DALL-E (Betker et al. 2023), and Stable Diffusion (Rombach et al. 2022; Podell et al. 2023). However, they often encounter challenges in generating images given complex and dense prompt descriptions, such as attribute binding and multiple objects (Huang et al. 2023).
@@
Figure 1: Architecture comparison between (a) LLM-guidance models (b) LLM-alignment models and (c) our proposed LLM4GEN.
With the emergence of powerful linguistic representations from Large Language Models (LLMs), there has been an increasing trend in leveraging LLMs to aid in T2I generation. Current methods mainly consist of two categories: LLM-guidance models (Yang et al. 2024; Feng et al. 2022) and LLM-alignment models (Wu et al. 2023; Hu et al. 2024; sd3; Zhao et al. 2024). LLM-guidance models harness the reasoning capability of LLMs and the Layout model to generate controllable images, as illustrated in Fig. 1 (a). However, these methods require separating LLMs from external models, resulting in redundancy in both inference time and the overall framework. While LLM-alignment models utilize LLMs to exploit the representational capacity, they demand substantial training data to align LLM representations with the diffusion model, as shown in Fig. 1 (b).
@@
To address aforementioned challenges, we propose LLM4GEN, a novel framework that implicitly leverages the powerful semantic representations of LLMs to enhance the original text encoder for T2I GENeration, as illustrated in Fig. 1 (c). Specifically, we design an efficient Cross-Adapter Module (CAM) to implicitly integrate the semantic representation of LLMs with original text encoders that have limited representational capabilities, such as CLIP text encoder (Radford et al. 2021). We apply cross-attention to the representations of both encoder-only LLMs (e.g., Llama (Zhang et al. 2024)) and decoder-only LLMs (e.g., T5 (Raffel et al. 2020)) alongside CLIP text embeddings, and then concatenate the fused embedding with the original CLIP text embedding. This CAM module significantly enhances the performance of T2I diffusion models while preserving the original text encoder representations, thereby reducing the need for extensive training data. Additionally, we introduce an entity-guidance regularization loss that penalizes mismatches between the activation maps of entities and their corresponding attributes in the text, improving the model’s ability to accurately comprehend and represent the main subjects in the generated images. As evidenced in Fig. 2, our proposed method exhibits strong performance in T2I generation.
Figure 2: Image generation using concise and dense prompts, with colored text highlighting key entities or attributes(Zoom in for details).
@@
To comprehensively assess the image generation capabilities of T2I models, we develop a comprehensive benchmark named DensePrompts, an extension of T2I-CompBench (Huang et al. 2023), which incorporates over 7,000 compositional prompts. The construction of this benchmark involves leveraging LLMs for complex text descriptions, followed by manual refinement. Results from performance metrics and human evaluations consistently demonstrate that LLM4GEN’s representational capability surpasses other existing methods.
@@
  • We propose a novel framework that leverages the powerful representational capabilities of LLMs to assist in text-to-image (T2I) generation. Specifically, we design a Cross-Adapter to integrate LLM representations and introduce an entity-guidance regularization loss to enhance semantic understanding.
  • To assess performance with long-text prompts, we introduce DensePrompts, a benchmark designed to evaluate both aesthetic quality and image-text alignment.
@@
  • Our designed LLM4GEN can be seamlessly integrated into existing diffusion models like SD1.5 (Rombach et al. 2022) and SDXL (Podell et al. 2023). Experiments show that LLM4GEN exhibits superior performance in sample quality, image-text alignment, and human evaluation compared with existing state-of-the-art models.
Related Work
Large Language Models
@@
Large language models (LLMs) (Chang et al. 2023) have shown powerful generalization ability in various NLP tasks. Recent LLMs, e.g., GPTs (Brown et al. 2020), LLaMA (Touvron et al. 2023), OPT (Zhang et al. 2022), PaLM (Chowdhery et al. 2022) are all equipped with billions of parameters, enabling the intriguing capability for in-context learning and demonstrating excellent zero-shot performance across various tasks. Certain Multi-modal LLMs (MLLMs) (Achiam et al. 2023; Zhu et al. 2023; Bai et al. 2023) have integrated visual and audio modalities, enhancing intelligent interactions with the help of LLMs. (Pang et al. 2024) shows that the frozen LLMs can further integrate visual understanding. Recent works (Lian et al. 2024; Sun et al. 2024; Yang et al. 2024) use LLMs to create improved text prompts or bounding box layouts for high-quality text-to-image generation. However, these existing works only consider LLMs as simple condition generators, e.g., text prompts or layout planning. In this paper, we harness the representation capabilities of LLMs to enhance text-to-image generation, emphasizing their significant representational power beyond simple text output.
@@
Text-to-Image Diffusion Models Text-to-image generation aims to create images with given prompts. Diffusion models (Song and Ermon 2019, 2020; Song et al. 2020) have demonstrated remarkable performance in image generation. These models use added Gaussian noise for a forward process and can generate diverse, high-quality images through an inverse process from random Gaussian noise. GLIDE (Nichol et al. 2022) utilizes CLIP (Radford et al. 2021) text encoder to enhance the image-text alignment. Latent Diffusion Models (LDMs) (Rombach et al. 2022) transfer the diffusion process from pixel to latent space. Recent models such as SD-XL (Podell et al. 2023), DALL-E 3 (Betker et al. 2023), and Dreambooth (Ruiz et al. 2023) have significantly enhanced image quality and text-image alignment using various perspectives, such as training strategies and scaling training data. Despite these notable advancements, generating high-fidelity images aligned with complex textual prompts remains challenging. In this paper, we propose LLM4GEN, which leverages the robust representation capabilities of LLMs to facilitate image generation from textual descriptions.
Figure 3: The overview of LLM4GEN. (a) Framework. (b) Cross-Adapter Module.
Methodology
LLM4GEN
Framework
The proposed LLM4GEN, which contains a Cross-Adapter Module (CAM) and the UNet, is illustrated in Fig. 3 (a). In this paper, we explore stable diffusion (Rombach et al. 2022; Podell et al. 2023) as the base text-to-image diffusion model, and the vanilla text encoder is from CLIP (Radford et al. 2021). LLM4GEN leverages the strong capability of LLMs to assist in text-to-image generation. The CAM extracts the representation of a given prompt via the combination of LLM and CLIP text encoder. The fused text embedding is enhanced by leveraging the pre-trained knowledge of LLMs through the simple yet effective CAM. By feeding the fused text embedding, LLM4GEN iteratively denoises the latent vectors with the UNet and decodes the final vector into an image with the VAE.
Cross-Adapter Module
$Q=W_{q}(c_{l}),K=W_{k}(c_{t}),V=W_{v}(c_{t})$ (1)
$c_{l}^{\prime}=\operatorname{CrossAttention}(Q,K,V)=\operatorname{softmax}\left(Q\cdot K^{T}\right)\cdot V$ (2)
$x=\operatorname{CA}(x,\operatorname{Concat}(\lambda\cdot c_{l}^{\prime},c_{t}))=\lambda\cdot\operatorname{CA}(x,c_{l})+\operatorname{CA}(x,c_{t})$ (3)
where $\operatorname{Concat}$ denotes concatenation in the sequence dimension, and $\lambda$ is the balance factor, $x$ denotes the latent noise, $\operatorname{CA}$ is the cross-attention module within the UNet module. Overall, our designed Cross-Adapter Module implicitly facilitates the strong representation of LLMs with a residual fusion manner, without utilizing extensive training data and resources to condition the latent vectors on text embeddings. Notably, our LLM4GEN is compatible with both decoder-only and encoder-only LLMs and we evaluate on Llama-2 7B/13B (Touvron et al. 2023) and T5-XL (Brown et al. 2020) in further experiments.
Figure 4: Statistic of DensePrompts benchmark compared with other benchmarks.
Entity-Guidance Regularization Loss
Current text-to-image generation models often encounter confusion and omissions when generating multiple entities. We utilize a parser to analyze the prompt $\mathcal{P}$, extracting a set of attribute-entity pairs $\mathcal{S}=\sum_{i=1}^{N}\{a_{i},e_{i}\}$, where $e_{i}$ and $a_{i}$ represent the entity name and its corresponding attribute, respectively, and $N$ denotes the number of parsed pairs. Subsequently, we can calculate the active similarity map as:
$\mathcal{A}_{i}=\operatorname{softmax}\left(\frac{QK_{i}^{T}}{\sqrt{d}}\right)$ (4)
where the query $Q$ is derived from the latent representation, the key $K_{i}$ is derived from the token embedding of $p$, and $d$ is the latent dimension. $\mathcal{A}_{a}$ and $\mathcal{A}_{o}$ indicate the similarity maps for the attribute word and the entity, respectively. Subsequently, we impose a penalty on these similarity maps on all UNet layers as:
$\mathcal{L}_{reg}=\frac{1}{N\cdot L}\sum_{i=1}^{N}\sum_{l=1}^{L}||\mathcal{A}_{a}^{i-l}-\mathcal{A}_{o}^{i-l}||^{2}$ (5)
where |||||$\cdot$||| | ⋅ | | represents the L2 distance, $L$ is the layer numbers.
Overall, based on the framework described above, the training loss of LLM4GEN is formulated as:
@@
$\mathcal{L}=\mathbb{E}_{\varepsilon(x),\epsilon\sim\mathcal{N}(0,1),t}[\parallel\epsilon-\epsilon_{\theta}(z_{t},t)\parallel]_{2}^{2}]+\alpha\cdot\mathcal{L}_{reg}$ (6)
@@
DensePrompts Benchmark
@@
A comprehensive benchmark is crucial for evaluating the image-text alignment of generated images. Current benchmarks, e.g., MSCOCO (Lin et al. 2014) and T2I-CompBench (Huang et al. 2023), primarily consist of concise textual descriptions, are not comprehensive enough to describe a diverse range of objects. Thus, we introduce a new comprehensive and complicated benchmark called DensePrompts, comprising lengthy textual descriptions.
@@
Initially, we collect 100 images from the Internet, comprising 50 real and 50 generated images, each with intricate details. Leveraging the robust image comprehension capabilities of GPT-4V (OpenAI 2023), we utilize it to provide detailed descriptions for these 100 images, encompassing object attributes and their relationships, thereby generating comprehensive prompts abundant in semantic details. We employ GPT-4 (Achiam et al. 2023) to produce massive long texts based on generated prompts mentioned above. DensePrompts provides more than 7,000 extensive prompts whose average word length is more than 40. Word statistics of DensePrompts are outlined in Fig. 4. To assess the performance, DensePrompts benchmark incorporates CLIP Score (Radford et al. 2021) and Aesthetic Score (Schuhmann ). Combining our proposed DensePrompts with T2I-CompBench, we establish a comprehensive evaluation in text-to-image generation.
Experiments
Experimental Details
Framework and Implementation Details
In this paper, we explore LLM4GEN based on SD1.5 and SDXL, denoted as LLM4GENSD1.5 and LLM4GENSDXL. We utilize T5-XL and CLIP text encoder (CLIP ViT-L/14) as the text tower. The sequence length of the LLMs is set to $128$. We use 10M text-image pairs collected from LAION-2B(Schuhmann et al. 2021) and Internet. Training is conducted on 8NVIDIA A100 GPUs with the learning rates of 2e-5 and 1e-5 for LLM4GENSD1.5 and LLM4GENSDXL, respectively. The batch size is set to 256 and 128. The training steps are set to 20k and 40k. Additionally, we further train LLM4GENSDXL using 2M high-quality data with 1024 resolution. During inference, we utilize DDIM sampler (Song, Meng, and Ermon 2020) for sampling with 50 steps and the classifier free guidance scale to 7.5.
Evaluation Benchmarks
We comprehensively evaluate proposed LLM4GEN via four primary benchmarks, e.g. MSCOCO (Lin et al. 2014), T2I-CompBench (Huang et al. 2023), our proposed DensePrompts benchmark, and User Study.
@@
Composable Diffusion (Liu et al. 2022) 40.63 32.99 36.45 8.00 29.80 28.98
Structured Diffusion (Feng et al. 2022) 49.90 42.18 49.00 13.86 31.11 33.55
Attn-Exct v2 (Chefer et al. 2023) 64.00 45.17 59.63 14.55 31.09 34.01
GORS (Huang et al. 2023) 66.03 47.85 62.87 18.15 31.93 33.28
DALL-E 2 (Ramesh et al. 2022) 57.50 54.64 63.74 12.83 30.43 36.96
PixArt-$\alpha$ (Chen et al. 2023b) 68.86 55.82 70.44 20.82 31.79 41.17
SUR-Adapter* (Zhong et al. 2023) 38.93 36.73 39.21 13.21 30.81 31.43
ELLASDXL (Hu et al. 2024) 72.60 56.34 66.86 22.14 30.69 -
SD1.5 (Rombach et al. 2022) 37.65 35.76 41.56 12.46 30.79 30.80
LLM4GENSD1.5 (Ours) 47.34 45.35 55.39 14.79 31.00 33.97
$\Delta$ (Margin) +9.69 +9.59 +13.83 +2.33 +0.21 +3.17
SDXL (Podell et al. 2023) 63.69 54.08 56.37 20.32 31.10 40.91
LLM4GENSDXL (Ours) 76.59 59.24 70.86 24.12 32.10 42.19
$\Delta$ (Margin) +12.90 +5.16 +14.49 +3.80 +1.00 +1.28
Table 1: Evaluation results (%) on T2I-CompBench (Huang et al. 2023). The higher is better, and the best results are highlighted in bold. * denotes that we calculate the metrics using the official weights and code provided.